OpenStudy (anonymous):

Simplify the expression (10^4 x 2^4) (10^3 x 2^3) / (10 x 2)^2

OpenStudy (anonymous):

\[\frac{ (10^4 x 2^4) (10^3 x 2^3) }{ (10 x 2)^2 }\]

jhonyy9 (jhonyy9):

10^4 *2^4 *10^3 *2^3 10^4 *2^4 *10^3 *2^3 ---------------------- = --------------------- = (10*2)^2 10^2 *2^2 10^2 *10^2 *2^2 *2^2 *10^3 *2^3 = -------------------------------- = 2^2 *2^2 *10^3 *2^3 = 2^4 *2^3 *10^3 = 10^2 *2^2 = 2^(4+3) *10^3 = 2^7 *10^3 - so step by step for more easy understandably - ok. ?